SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (Aug 9, 2005)

Attendees: Rainer Bartoldus, Kris Chaisanguanthum, Al Eisner, Selina Li, Brian Lindquist, Dong Su.

Operations

  • Selina will be giving the monthly Trigger status report tomorrow after the detector operation. A copy of the talk will be linked from here when it's available.

EMT

  • Dan will be testing the EMT prescaling from database at the test stand this Wednesday before deploying it at IR2.

DCT/DCZ

  • Jamie suggested to swap the ZPD boards to investigate the problem with a dip seen at phi ~60 degrees in the efficiency plot. See HN for more information. Two ZPD boards were swapped on August 4 at IR2, but no effect took place. The conclusion was that it's due to the DCH frontend electronics since both the new and old TSF occupancy plots show a dip at this region as well.

GLT

  • We tested the GLT FC2 version of the firmware at IR2 last Tuesday and Wednesday (Aug.2nd and 3rd). There were synchrequest errors seen by DAQ with this version of the firmware which wasn't observed at the test stand. Finally, Su Dong was able to reproduce the error on both platform 27 and 30 by probing the monitor hook on the clock. It was close on edge in timing that cause the damage errors. Work is under way to get another version working.

L3

  • The L3 tracking bug fix will be in the nex online release for sure. This will solve the problem with the infrequent L3 node trim.

Data Quality and Monitoring

  • Mike Kelsey requested us to include number of hits per track for L3 monitoring for diagnostic purpose.
  • We discussed briefly about offline luminosity compared to L3 lumi. Al suggested to add some monitoring plots such as lumi vs bhabha accepts, and to create a stripchart for those quantities in the trigger operation page first. Su Dong pointed out that looking at "bhabha accept OR EMC bhabha" will be more reliable.
  • Rainer suggested to put the boost quantity to help monitor beam energy as well since it's very sensitive to change in energy for one beam.
  • Kris is looking at the ZPD monitoring code and found that there are some BLT/ZPD digit modules that mixed with that of the BLT/PTD's. This dependency needs to be removed in case we need to take out the old DCT system from production.
  • The L1T crate damage check has updated instruction in JAS now.
  • In checking the reprocessed Run 2 data, Kris found out that there were about 60 runs (e.g. run 23624) taken around October 2001 having a dip in the TSF efficiency plot. The dip obviously showed up in the efficiency plots for Ap and A. However, we couldn't find a record of this problem in the elog, trigger history record or HN. This is a potential problem that DCH and tracking groups should take a look at their relevant plots.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.